# The adjacent Hindman’s theorem and the $\mathbb{Z}$-Ramsey theorem

### v1

@@ 2.1. Relation with the Adjacent Hindman’s Theorem
Now, to prove $\mathnormal{\mathsf{AHT}}^{d}\leq_{u}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}$, the witnessing Turing functionals will be $\Phi_{2},\Psi_{2}$ defined as follows: for $d:\mathbb{N}^{d}\longrightarrow k$ we let $\Phi_{2}^{d}:[\mathbb{N}]^{d+1}\longrightarrow k$ be given by $\Phi_{2}^{d}(x_{0},\ldots,x_{d})=d(x_{1}-x_{0},x_{2}-x_{1},\ldots,x_{d}-x_{d-1})$ (if $x_{1}<\ldots<x_{d+1}$); it is easy to see that $\Phi_{2}^{d}$ is a $\mathbb{Z}$-invariant colouring. For the definition of $\Psi_{2}$ we proceed as in the proof of theorem 5: given an infinite set $X$, first define $X^{\prime}=\{x_{n}\big{|}n\in\mathbb{N}\}$ such that $x_{0},x_{1}$ are the two smallest elements of $X$, and $x_{n+1}$ is the least element of $X$ so that $x_{n+1}-x_{n}>x_{n}-x_{n-1}$ for each $n\geq 1$. We then let $\Psi_{2}^{X}=\{x_{n+1}-x_{n}\big{|}n\in\mathbb{N}\}$. Note that, if $y_{j}=x_{j+1}-x_{j}$, then an element $\vec{y}\in\operatorname{AFS}^{d}(\Psi_{2}^{X})$ is of the form
@@
so that $d(\vec{y})=\Phi_{2}^{d}(x_{0},\ldots,x_{d})$ and hence, if $[X]^{n+1}$ is $\Phi_{2}^{d}$-monochromatic this implies that $\operatorname{AFS}(\Psi_{2}^{X})$ is $d$-monochromatic. ∎
As clearly as always, this leads also to the following.
@@
For each $d$, the statements -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}$ and $\mathnormal{\mathsf{AHT}}^{d}$ are equivalent over $\mathnormal{\mathsf{RCA}_{0}}$.
@@ 3.1. The apartness condition
In Carlucci’s paper [2] (where the original formulation of the $\mathnormal{\mathsf{AHT}}$ can be found), a version of the $\mathnormal{\mathsf{AHT}}$ is considered where one requires an apartness condition on the generators of the monochromatic set. In order to understand this requirement, recall that given an $x\in\mathbb{N}$ one defines $\lambda(x)$ to be the unique $n$ such that $2^{n}\mid x$ but $2^{n+1}\nmid x$; on the other hand, one defines $\mu(x)=\lfloor\log_{2}(x)\rfloor$. Pictorially, it is easier to remember that, when expressing $x$ in binary notation, $\lambda(x)$ provides the position of the first non-zero digit and $\mu(x)$ is the position of the last (non-zero) digit (first and last meaning from lower to higher powers of $2$ —i.e., when reading the number from right to left); even more intuitively, one can simply think of the natural number $n$ as a finite subset of $\mathbb{N}$ (by letting the binary expansion of $n$ represent the characteristic function of such finite subset) and then $\lambda(n),\mu(n)$ are simply the minimum and maximum elements of this finite set.
@@
The reader familiar with the definition of a block sequence on $[\mathbb{N}]^{<\aleph_{0}}$ will recognize that a sequence of natural numbers satisfies the apartness condition precisely when the finite sets corresponding to the elements of the sequence (as described in the previous paragraph) form a block sequence. Carlucci’s adjacent Hindman’s theorem with apartness is the principle stating the same as $\mathnormal{\mathsf{AHT}}$ but requiring that the sequence whose set of adjacent finite sums is monochromatic satisfy the apartness condition; we will henceforth denote this principle with the symbol $\mathnormal{\mathsf{AHT}}_{a}$. We similarly define the higher-dimensional versions $\mathnormal{\mathsf{AHT}}_{a}^{d}$ for each $d$.
@@
    A sequence $\langle x_{n}\big{|}n\in\mathbb{N}\rangle$ is said to satisfy the separation condition if, for every $n\in\mathbb{N}$, we have $\mu(x_{n}-x_{n-1})<\lambda(x_{n+1}-x_{n})$.
@@
    Given an $n\in\mathbb{N}$, the $\mathbb{Z}$-Ramsey theorem for $n$-tuples with separation is the statement that every $\mathbb{Z}$-invariant colouring of $n$-tuples admits an infinite monochromatic set that, when enumerated increasingly, yields a sequence satisfying the separation condition. We denote this principle with the symbol -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{n}$.
Clearly, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{n}$ implies -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ in $\mathnormal{\mathsf{RCA}_{0}}$ or, in terms of uniform reducibility, we have -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}\leq\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{n}$ (intuitively, requiring the separation condition in the monochromatic set makes the corresponding statement stronger). Looking carefully (whether literally as Computability statements, or as proofs in $\mathnormal{\mathsf{RCA}_{0}}$) at the proofs from Theorem 8 (as described two paragraphs above but now in this more general context), yields the following.
@@
    -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{d+1}\leq_{u}\mathnormal{\mathsf{AHT}}_{a}^{d}$ and $\mathnormal{\mathsf{AHT}}_{a}^{d}\leq_{u}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{d+1}$,
@@
    for each $d$, the statements -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{d+1}$ and $\mathnormal{\mathsf{AHT}}_{a}^{d}$ are equivalent over $\mathnormal{\mathsf{RCA}_{0}}$.
For a relationship between the versions of -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}$ with or without the separation condition (or, equivalently, between the versions of $\mathnormal{\mathsf{AHT}}$ with or without apartness), Carlucci [2] proves that $\mathnormal{\mathsf{RT}}^{2}$ implies $\mathnormal{\mathsf{AHT}}_{a}$. Using essentially the same reasoning, we prove the higher-dimensional general version of this result.
@@
For each $n\in\mathbb{N}$, we have $\mathnormal{\mathsf{AHT}}_{a}^{n}\leq_{u}\mathnormal{\mathsf{RT}}^{n+1}$.
@@
Then $[Y]^{n+1}$ is $\Phi^{c}$-monochromatic if and only if $\operatorname{AFS}^{n}(\Psi^{Y})$ is $c$-monochromatic. ∎
This way, we obtain the following diagram, where the arrows can mean either implication under $\mathnormal{\mathsf{RCA}_{0}}$, or uniform reducibility (of the item to the right of the arrow to the item to the left of it)
[…]
Towards the right-extreme of this diagram (which extends infinitely to the left), we find the following configuration:
[…]
This is very informative from the perspective of uniform reducibility. From the perspective of Reverse Mathematics, on the other hand, a huge portion of the diagram collapses, since it is known that $\mathnormal{\mathsf{RT}}^{3}$ already implies $\mathnormal{\mathsf{ACA}_{0}}$. Therefore, the first diagram consists of equivalent statements, and the only question remaining should be regarding the second diagram. The next (and last) section of the paper addresses this question.
@@
In this section, we show that $\mathnormal{\mathsf{AHT}}^{2}$ implies $\mathnormal{\mathsf{ACA}_{0}}$ over $\mathnormal{\mathsf{RCA}_{0}}$. Thus, the leftmost five nodes of the last shown diagram collapse, all of them being equivalent, and really the only point of separation happens in the arrow connecting -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}$ to $\mathnormal{\mathsf{RT}}^{2}$. The following result, stated in both the language of Computability Theory and of Reverse Mathematics, is proved with much the same ideas as the classical proof that $\mathnormal{\mathsf{RT}}^{3}$ implies $\mathnormal{\mathsf{ACA}_{0}}$.
@@
Consider the colouring $c:\mathbb{N}^{2}\longrightarrow 2\times 2$ given by letting $c(x,y)=(i,j)$, where $i$ equals $1$ if and only if $\lambda(x)<\lambda(y)$, and $j$ equals $1$ if and only if $(0^{\prime}\cap\lambda(x))[\mu(x)]=(0^{\prime}\cap\lambda(x))[\mu(y)]$. Suppose that $\vec{x}=(x_{n}\big{|}n\in\mathbb{N})$ is a sequence such that $\operatorname{AFS}^{2}(\vec{x})$ is $c$-monochromatic. Let $n_{0}$ be such that $\lambda(x_{n_{0}})$ is as small as possible. Comparing $\lambda(x_{n_{0}})$ with $\lambda(x_{n_{0}+1})$ and $\lambda(x_{n_{0}+2})$, we see that either $\lambda(x_{n_{0}})<\lambda(x_{n_{0}+1})$, or $\lambda(x_{n_{0}+1})<\lambda(x_{n_{0}+2})$, or $\lambda(x_{n_{0}})=\lambda(x_{n_{0}+1})=\lambda(x_{n_{0}+2})$ which implies $\lambda(x_{n_{0}})<\lambda(x_{n_{0}+1}+x_{n_{0}+2})$. In either case, we have obtained an element of $\operatorname{AFS}^{2}(\vec{x})$ whose colour is of the form $(1,j)$; hence all elements of $\operatorname{AFS}^{2}(\vec{x})$ have the same colour and so $\lambda$ is strictly increasing on the sequence $\vec{x}$. In particular, this implies that $\lambda(x_{n}+x_{n+1}+\cdots+x_{n+l})=\lambda(x_{n})$ for all $n,l$. Therefore, for any $n$ it is possible to take an $l$ large enough that $(0^{\prime}\cap\lambda(x_{n}))[\mu(x)]=0^{\prime}\cap\lambda(x_{n})$ for $x=x_{n}+x_{n+1}+\cdots+x_{n+l}$; then any element of $\operatorname{AFS}^{2}(\vec{x})$ having $x$ as its first coordinate will have colour $(1,1)$. This implies that for every $n\in\mathbb{N}$ we have $(0^{\prime}\cap\lambda(x_{n}))[\mu(x_{n})]=0^{\prime}\cap\lambda(x_{n})$ —otherwise, taking any $t$ large enough that $(0^{\prime}\cap\lambda(x_{n}))[\mu(y)]=0^{\prime}\cap\lambda(x_{n})$ for $y=x_{n+1}+\ldots+x_{n+t}$, we would obtain the element $(x_{n},y)\in\operatorname{AFS}^{2}(\vec{x})$ with colour $(1,0)$—. In other words, the sequence $\vec{x_{n}}$ is able to compute $0^{\prime}$ (given any $m\in\mathbb{N}$, take an $n$ large enough that $\lambda(x_{n})>m$ and check whether $m$ belongs to $(0^{\prime}\cap\lambda(x_{n}))[\mu(x_{n})]$). ∎
${\mathnormal{\mathsf{ACA}_{0}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}_{s}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}}$${\mathnormal{\mathsf{AHT}}_{a}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}}$pigeonhole
@@
Figure 1. Diagram of implications over $\mathnormal{\mathsf{RCA}_{0}}$.
@@
We summarize the results of the paper, from the perspective of Reverse Mathematics, by presenting the diagram in Figure 1. Whether any of the remaining arrows (except for the one connecting $\mathnormal{\mathsf{ACA}_{0}}$ to $\mathnormal{\mathsf{RT}}^{2}$) is reversible remains an open question. For example, it is conceivable that $\mathnormal{\mathsf{AHT}}$ already implies $\mathnormal{\mathsf{AHT}}_{a}$, although some ideas that have been used for similar results (e.g. [6, Lemma 1]) do not seem to work in this context because the requirement that the sums be adjacent, thus preventing us to skip summands, poses a strong restriction. Similar difficulties arise when one attempts to obtain better lower bounds for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$: Carlucci [2, Prop. 3] has found that the Increasing Polarized Ramsey’s Theorem for pairs [^2] That is, the statement that for every colouring $c:[\mathbb{N}]^{2}\longrightarrow k$ there are two infinite sets $X,Y$ such that all pairs of the form $\{x,y\}$ with $x\in X$, $y\in Y$, and $x<y$, have the same colour., denoted $\mathsf{IPT}^{2}$, is a lower bound for $\mathnormal{\mathsf{AHT}}_{a}$; the apartness condition plays a crucial rôle in Carlucci’s proof, and we have been unable to successfully find a suitable lower bound for $\mathnormal{\mathsf{AHT}}$ without this condition. We do have a conjecture but state it in the form of a question below. A final consideration arises from thinking about versions of the principles we have stated for a bounded number of colours. For example, one could consider, e.g., the principle $\mathnormal{\mathsf{AHT}}_{k}$ which states the same as $\mathnormal{\mathsf{AHT}}$ but only for colourings with at most $k$ colours. Note that, from this viewpoint, Theorem 14 really proves that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}_{4}$ implies $\mathnormal{\mathsf{ACA}_{0}}$, but it is not clear whether that $4$ could be lowered even further. We finalize the paper by stating explicitly these questions.

### v2

@@ 2.1. Relation with the Adjacent Hindman’s Theorem
Now, to prove $\mathnormal{\mathsf{AHT}}^{d}\mathop{\leq_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}$, the witnessing Turing functionals will be $\Phi_{2},\Psi_{2}$ defined as follows: for $d:\mathbb{N}^{d}\longrightarrow k$ we let $\Phi_{2}^{d}:[\mathbb{N}]^{d+1}\longrightarrow k$ be given by $\Phi_{2}^{d}(x_{0},\ldots,x_{d})=d(x_{1}-x_{0},x_{2}-x_{1},\ldots,x_{d}-x_{d-1})$ (if $x_{1}<\ldots<x_{d+1}$); it is easy to see that $\Phi_{2}^{d}$ is a $\mathbb{Z}$-invariant colouring. For the definition of $\Psi_{2}$ we proceed as in the proof of theorem 5: given an infinite set $X$, first define $X^{\prime}=\{x_{n}\big|n\in\mathbb{N}\}$ such that $x_{0},x_{1}$ are the two smallest elements of $X$, and $x_{n+1}$ is the least element of $X$ so that $x_{n+1}-x_{n}>x_{n}-x_{n-1}$ for each $n\geq 1$. We then let $\Psi_{2}^{X}=\{x_{n+1}-x_{n}\big|n\in\mathbb{N}\}$. Note that, if $y_{j}=x_{j+1}-x_{j}$, then an element $\vec{y}\in\operatorname{AFS}^{d}(\Psi_{2}^{X})$ is of the form
@@
so that $d(\vec{y})=\Phi_{2}^{d}(x_{0},\ldots,x_{d})$ and hence, if $X]^{n+1}$ is $\Phi_{2}^{d}$-monochromatic this implies that $\operatorname{AFS}(\Psi_{2}^{X})$ is $d$-monochromatic. ∎
Once again, by going carefully through the previous proof, and noticing that (the proof is simple enough that) no induction beyond $\mathnormal{\mathsf{I}\mathrm{\Sigma}_{1}^{0}}$ is used, one proves the following corollary.
@@
For each $d$, $\mathnormal{\mathsf{RCA}}_{0}\vdash\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\leftrightarrow\mathnormal{\mathsf{AHT}}^{d}$.
@@ 3.1. The apartness condition
In Carlucci’s paper [2] (where the original formulation of the $\mathnormal{\mathsf{AHT}}$ can be found), a version of the $\mathnormal{\mathsf{AHT}}$ is considered where one requires an apartness condition on the generators of the monochromatic set. In order to understand this requirement, recall that, after [1], for an $x\in\mathbb{N}$ we let $\lambda(x)$ be the smallest position, and we let $\mu(x)$ be the largest position, where a nonzero digit appears if $x$ is written in binary notation (when reading the number from right to left, i.e., from the least significant to the most significant digit). Equivalently, and even more intuitively, one can simply think of the natural number $n$ as a finite subset of $\mathbb{N}$ (by letting the binary expansion of $n$ represent the characteristic function of such finite subset) and then $\lambda(n)$ and $\mu(n)$ are simply the minimum and maximum elements of this finite set. There are other ways to describe these couple of functions (for example, $\lambda(x)$ is the unique $n$ such that $2^{n}\mid x$ but $2^{n+1}\nmid x$, and $\mu(x)=\lfloor\log_{2}(x)\rfloor$), but the description in terms of the binary representation is much easier to visualize. Using this tool, we are able to state the following definition (which is already implicit in [1], and more explicit in [2, Definition 3], but for sets; we simply render below the same definition but for sequences).
@@
The reader familiar with the definition of a block sequence on $\mathbb{N}]^{<\aleph_{0}}$ will recognize that a sequence of natural numbers satisfies the apartness condition precisely when the finite sets corresponding to the elements of the sequence (as described in the previous paragraph) form a block sequence. Carlucci’s adjacent Hindman’s theorem with apartness is the principle stating the same as $\mathnormal{\mathsf{AHT}}$ but requiring that the sequence whose set of adjacent finite sums is monochromatic satisfy the apartness condition; we will henceforth denote this principle with the symbol $\mathnormal{\mathrm{ap}\mathsf{AHT}}$. We similarly define the higher-dimensional versions $\mathnormal{\mathrm{ap}\mathsf{AHT}}^{d}$ for each $d$.
@@
    A set $X$ is said to satisfy the separation condition if, letting $\langle x_{n}\big|n\in\mathbb{N}\rangle$ be the increasing enumeration of $X$, we have $\mu(x_{n}-x_{n-1})<\lambda(x_{n+1}-x_{n})$ for every $n\in\mathbb{N}$.
@@
    Given an $n\in\mathbb{N}$, the $\mathbb{Z}$-Ramsey’s theorem for $n$-subsets with separation is the statement that every $\mathbb{Z}$-invariant colouring of $n$-subsets admits an infinite monochromatic set that, when enumerated increasingly, yields a sequence satisfying the separation condition. We denote this principle with the symbol $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{n}$.
Notice that the separation condition is not hereditary to subsets, i.e., if $Y\subseteq X$ then $Y$ does not necessarily satisfy the separation condition even if $X$ does; therefore, $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}$ is a priori a stronger statement than -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}$ (it cannot be obtained simply by applying -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}$ to a set that satisfies the separation condition; furthermore, it is not even clear that it can be obtained by first applying -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}$ and then thinning out the resulting monochromatic set). More concretely, $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ implies -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}$ in $\mathnormal{\mathsf{RCA}}_{0}$ or, in terms of Weihrauch reducibility, we have -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}\leq\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{n}$. Looking carefully (whether literally as Computability statements, or as proofs in $\mathnormal{\mathsf{RCA}}_{0}$) at the proofs from Theorem 8 (as described two paragraphs above but now in this more general context), yields the following.
@@
    $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}\mathop{\equiv_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathrm{ap}\mathsf{AHT}}^{d}$,
@@
    for each $d$, the statements $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{d+1}$ and $\mathnormal{\mathrm{ap}\mathsf{AHT}}^{d}$ are equivalent over $\mathnormal{\mathsf{RCA}}_{0}$.
For a relationship between the versions of -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}$ with or without the separation condition (or, equivalently, between the versions of $\mathnormal{\mathsf{AHT}}$ with or without apartness), Carlucci [2] proves that $\mathnormal{\mathsf{RT}}^{2}$ implies $\mathnormal{\mathrm{ap}\mathsf{AHT}}$. Using the same reasoning, we prove the higher-dimensional general version of this result.
@@
For each $n\in\mathbb{N}$, we have $\mathnormal{\mathrm{ap}\mathsf{AHT}}^{n}\mathop{\leq_{\mathrm{s}\mathsf{W}}}\mathnormal{\mathsf{RT}}^{n+1}$.
@@
Then $Y]^{n+1}$ is $\Phi^{c}$-monochromatic if and only if $\operatorname{AFS}^{n}(\Psi^{Y})$ is $c$-monochromatic. ∎
This way, we obtain the following diagram, where the arrows can mean either implication under $\mathnormal{\mathsf{RCA}}_{0}$, or Weihrauch reducibility (and we make no further claims as to whether the arrows are reversible or not; some of these collapse under $\mathnormal{\mathsf{RCA}}_{0}$ but not necessarily under Weihrauch reduction)
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n+1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots}\textstyle{\cdots}$${\mathnormal{\mathrm{ap}\mathsf{AHT}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}^{n}}$${\mathnormal{\mathrm{ap}\mathsf{AHT}}^{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}^{n-1}}\textstyle{\cdots}$
Towards the right-extreme of this diagram, we find the following configuration:
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\mathnormal{\mathsf{RT}}^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots}$${\mathnormal{\mathrm{ap}\mathsf{AHT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}^{2}}$${\mathnormal{\mathrm{ap}\mathsf{AHT}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}}$pigeonhole
(Other principles could be added to the diagram, e.g., the fact that $\mathnormal{\mathrm{ap}\mathsf{AHT}}$ implies $\mathsf{D}^{2}$ [2], but we only added the principles explicitly studied in this paper to the diagram.) This is very informative from the perspective of Weihrauch reducibility. From the perspective of Reverse Mathematics, on the other hand, a huge portion of the diagram collapses, since it is known that $\mathnormal{\mathsf{RT}}^{3}$ already implies $\mathnormal{\mathsf{ACA}}_{0}$. Therefore, the first diagram (up to $n=4$) consists of equivalent statements, and the only question remaining should be regarding the second diagram. The next (and last) section of the paper addresses this question.
@@
In this section, we show that $\mathnormal{\mathsf{AHT}}^{2}$ implies $\mathnormal{\mathsf{ACA}}_{0}$ over $\mathnormal{\mathsf{RCA}}_{0}$. Thus, the leftmost five nodes of the last shown diagram collapse, all of them being equivalent, and really the only point of separation happens in the arrow connecting -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}$ to $\mathnormal{\mathsf{RT}}^{2}$. The following result, stated in both the language of Computability Theory and of Reverse Mathematics, is proved with much the same ideas as the classical proof that $\mathnormal{\mathsf{RT}}^{3}$ implies $\mathnormal{\mathsf{ACA}}_{0}$ [^3] There are other examples of statements where dimension 2 suffices to prove $\mathnormal{\mathsf{ACA}}_{0}$, such as the Canonical Ramsey’s Theorem, the Regressive Function’s Theorem [13] or the Regressive Hindman’s Theorem [4]..
@@
We write explicitly the proof of (1), and only mention in passing that it readily formalizes to yield (2). Consider the colouring $c:\mathbb{N}^{2}\longrightarrow 2\times 2$ given by letting $c(x,y)=(i,j)$, where $i$ equals $1$ if and only if $\lambda(x)<\lambda(y)$, and $j$ equals $1$ if and only if $(0^{\prime}\cap\lambda(x))[\mu(x)]=(0^{\prime}\cap\lambda(x))[\mu(y)]$. Suppose that $\vec{x}=(x_{n}\big|n\in\mathbb{N})$ is a sequence such that $\operatorname{AFS}^{2}(\vec{x})$ is $c$-monochromatic. Let $n_{0}$ be such that $\lambda(x_{n_{0}})$ is as small as possible. Comparing $\lambda(x_{n_{0}})$ with $\lambda(x_{n_{0}+1})$ and $\lambda(x_{n_{0}+2})$, we see that either $\lambda(x_{n_{0}})<\lambda(x_{n_{0}+1})$, or $\lambda(x_{n_{0}+1})<\lambda(x_{n_{0}+2})$, or $\lambda(x_{n_{0}})=\lambda(x_{n_{0}+1})=\lambda(x_{n_{0}+2})$ which implies $\lambda(x_{n_{0}})<\lambda(x_{n_{0}+1}+x_{n_{0}+2})$. In either case, we have obtained an element of $\operatorname{AFS}^{2}(\vec{x})$ whose colour is of the form $(1,j)$; hence all elements of $\operatorname{AFS}^{2}(\vec{x})$ have the same colour and so $\lambda$ is strictly increasing on the sequence $\vec{x}$. In particular, this implies that $\lambda(x_{n}+x_{n+1}+\cdots+x_{n+l})=\lambda(x_{n})$ for all $n,l$. Therefore, for any $n$ it is possible to take an $l$ large enough that $(0^{\prime}\cap\lambda(x_{n}))[\mu(x)]=0^{\prime}\cap\lambda(x_{n})$ for $x=x_{n}+x_{n+1}+\cdots+x_{n+l}$; then any element of $\operatorname{AFS}^{2}(\vec{x})$ having $x$ as its first coordinate will have colour $(1,1)$. This implies that for every $n\in\mathbb{N}$ we have $(0^{\prime}\cap\lambda(x_{n}))[\mu(x_{n})]=0^{\prime}\cap\lambda(x_{n})$ —otherwise, taking any $t$ large enough that $(0^{\prime}\cap\lambda(x_{n}))[\mu(y)]=0^{\prime}\cap\lambda(x_{n})$ for $y=x_{n+1}+\ldots+x_{n+t}$, we would obtain the element $(x_{n},y)\in\operatorname{AFS}^{2}(\vec{x})$ with colour (1,0)(1,0)—. In other words, the sequence $\vec{x_{n}}$ is able to compute $0^{\prime}$ (given any $m\in\mathbb{N}$, take an $n$ large enough that $\lambda(x_{n})>m$ and check whether $m$ belongs to $(0^{\prime}\cap\lambda(x_{n}))[\mu(x_{n})]$). ∎
${\mathnormal{\mathsf{ACA}}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{RT}}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$-${\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}}$${\mathnormal{\mathrm{ap}\mathsf{AHT}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathnormal{\mathsf{AHT}}}$pigeonhole
@@
Figure 1. Diagram of implications over $\mathnormal{\mathsf{RCA}}_{0}$. The double arrow indicates that we know the implication is not reversible.
@@
We summarize the results of the paper, from the perspective of Reverse Mathematics, by presenting the diagram in Figure 1. Whether any of the remaining arrows (except for the one connecting $\mathnormal{\mathsf{ACA}}_{0}$ to $\mathnormal{\mathsf{RT}}^{2}$) is reversible remains an open question. For example, it is conceivable that $\mathnormal{\mathsf{AHT}}$ already implies $\mathnormal{\mathrm{ap}\mathsf{AHT}}$, although some ideas that have been used for similar results (e.g. [10, Lemma 1]) do not seem to work in this context because the requirement that the sums be adjacent, thus preventing us to skip summands, poses a strong restriction. Similar difficulties arise when one attempts to obtain better lower bounds for -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$: Carlucci [2, Prop. 3] has found that the Increasing Polarized Ramsey’s Theorem for pairs [^4] That is, the statement that for every finite colouring $c:[\mathbb{N}]^{2}\longrightarrow k$ there are two infinite sets $X,Y$ such that all pairs of the form $\{x,y\}$ with $x\in X$, $y\in Y$, and $x<y$, have the same colour., denoted $\mathsf{IPT}^{2}$, is a lower bound for $\mathnormal{\mathrm{ap}\mathsf{AHT}}$; the apartness condition plays a crucial rôle in Carlucci’s proof, and we have been unable to successfully find a suitable lower bound for $\mathnormal{\mathsf{AHT}}$ without this condition [^5] It is conceivable that some of the ideas from [5] could be useful to obtain such lower bound results without separation/apartness. We are grateful to one of the anonymous referees for this suggestion.. We do have a conjecture but state it in the form of a question below. A final consideration arises from thinking about versions of the principles we have stated for a bounded number of colours. For example, one could consider, e.g., the principle $\mathnormal{\mathsf{AHT}}_{k}$ which states the same as $\mathnormal{\mathsf{AHT}}$ but only for colourings with at most $k$ colours. Note that, from this viewpoint, Theorem 14 really proves that -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{3}_{4}$ implies $\mathnormal{\mathsf{ACA}}_{0}$, but it is not clear whether that $4$ could be lowered even further. We finalize the paper by stating explicitly these questions.